OpenStudy (xapproachesinfinity):

Curious if this is the right way looking for the convergence of \[\sum \sin (\frac{4n^2+5}{n^4+6})\]

OpenStudy (xapproachesinfinity):

this is what i did \[\frac{4n^2+5}{n^4+6}<\frac{4}{n^4}\] then \[\sin (\frac{4n^2+5}{n^4+6})<\sin (\frac{4}{n^4})\] the series on the right converges using p series test

OpenStudy (xapproachesinfinity):

I'm using the fact that sin is continuous therefore \[\sum \sin (4/n^4)=sin(\sum 4/n^4)\]

ganeshie8 (ganeshie8):

Your conclusion is correct, but reasoning isn't.. sin(x) is not a strictly increasing function

OpenStudy (xapproachesinfinity):

oh one little mistake the right hand of the inequality should have been 4/n^2

ganeshie8 (ganeshie8):

you may use this fact for \(x\gt 0\) instead : \[\sin(x) \lt x\]

OpenStudy (xapproachesinfinity):

hmm i see, just look for the convergence of what's inside using that fact

OpenStudy (xapproachesinfinity):

and conclude the convergent by comparison test

OpenStudy (xapproachesinfinity):

gotcha :) thanks

ganeshie8 (ganeshie8):

Yes\[\sin \frac{4n^2+5}{n^4+6} \lt\frac{4n^2+5}{n^4+6}<\frac{4n^2+5}{n^4} \] need to massage further..

OpenStudy (xapproachesinfinity):

that would be less than 4/n^2 yes?

OpenStudy (xapproachesinfinity):

which converges by p series

ganeshie8 (ganeshie8):

i don't think thats obvious..

ganeshie8 (ganeshie8):

how is \[\dfrac{4n^2+5}{n^4}\lt \dfrac{4}{n^2}\] ?

OpenStudy (xapproachesinfinity):

yeah it really is not hehe. we use limit comparison test

ganeshie8 (ganeshie8):

thats a good idea !!

OpenStudy (xapproachesinfinity):

but what sequence are we gonna use! i still have some subtlety with this test 4/n^2

OpenStudy (xapproachesinfinity):

lim 4/n^2/ (4n^2+6)/n^4 how do we decide the right sequence to do the limit comparison test

OpenStudy (xapproachesinfinity):

the book just mentions that the denominator behaves like a know sequence then they conclude the convergence with the limit if it is finite or not! I'm using the same technique, but it is not a guarantee for some cases

OpenStudy (xapproachesinfinity):

known*

OpenStudy (xapproachesinfinity):

gotta go, thanks for the help :)

ganeshie8 (ganeshie8):

that should work nicely right \[a_n:=\dfrac{4n^2+5}{n^4} \\~\\b_n:= \dfrac{1}{n^2}\] \[\lim\limits_{n\to\infty}~\dfrac{a_n}{b_n}=\lim\limits_{n\to\infty}~\dfrac{4n^2+5}{n^2}=4\] Since \(4\) is finite and the series \(\sum\limits_{n=1}^{\infty}b_n\) converges, the series \(\sum\limits_{n=1}^{\infty}a_n\) also converges.